The second angle of a triangle is 4 times as large as the first. The third angle is 110 more than the sum of the other two angles. Find the measure of the second angle.
Answer by nerdybill(7384)

The second angle of a triangle is 4 times as large as the first. The third angle is 110 more than the sum of the other two angles. Find the measure of the second angle.
.
Let x = first angle
then
4x = second angle
x+4x+110 = third angle
.
x + 4x + x+4x+110 = 180
10x+110 = 180
10x = 70
x = 7 degrees
.
Second angle:
4x = 4(7) = 28 degrees